OpenStudy (anonymous):

transform the equation into the corresponding polar coordinate equations. a.) x^2-2y^2=4

OpenStudy (anonymous):

\[x = rcos \theta,~~ y = r \sin \theta \]

OpenStudy (raden_zaikaria):